# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

### v1

@@ 4.3.2 Analysis of infinitesimal Hilbertianity
where in the last two steps, we first used the fact that $02|\nabla_{g}f_{1,n}|_{g}^{2}+2|\nabla_{g}f_{2,n}|_{g}^{2}-|\nabla_{g}f_{3,n}|_{g}^{2}-|\nabla_{g}f_{4,n}|_{g}^{2}=$ 0 for all $n\in\mathbb{N}$ and then applied (4.98). We have that $|Df_{i,1}|^{2}_{w}\in L^{\infty}$ and hence the last integral in the above equation is finite, so as $\mathrm{vol}_{g}(S\times(0,1))>0$, we get that
$\lim_{n\to\infty}2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})\infty.$ (4.99)
In particular, there exists an $n\in\mathbb{N}$ such that $2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})\neq 0$. That shows:
Proposition 4.12.
This proves Theorem 2.
@@
For a manifold with a $W^{1,p}_{\rm loc}$-Riemannian metric such that $p>d$, the Sobolev embedding theorem together with the results from [28] immediately give us that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian. Theorem 2 shows that $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ for $p<d-1$ is not sufficient for infinitesimal Hilbertianity. This works towards answering Question 3 leaving only the interval of $p\in[d-1,d]$, where it is unknown whether $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ is sufficient to get infinitesimal Hilbertianity.

### v3

@@ 4.3.2 Analysis of infinitesimal Hilbertianity
where in the last two steps, we first used the fact that $2|\nabla_{g}f_{1,n}|_{g}^{2}+2|\nabla_{g}f_{2,n}|_{g}^{2}-|\nabla_{g}f_{3,n}|_{g}^{2}-|\nabla_{g}f_{4,n}|_{g}^{2}=0$ for all $n\in\mathbb{N}$ and then applied (4.105). We have that $|Df_{i,1}|^{2}_{w}\in L^{\infty}$ and hence the last integral in the above equation is finite, so as $\mathrm{vol}_{g}(S\times(0,1))>0$, we get that
$\lim_{n\to\infty}2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})=\infty.$ (4.106)
In particular, there exists an $n\in\mathbb{N}$ such that $2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})\neq 0$. ∎
This proves Theorem 2, taking into account that for $p^{\prime}\in[1,d-1]$, one has $W^{1,p^{\prime}}_{\rm loc}\subset W^{1,d-1}_{\rm loc}$.
@@
For a manifold with a $W^{1,p}_{\rm loc}$-Riemannian metric such that $p>d$, the Sobolev embedding theorem together with the results from [30] immediately give us that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian. Theorem 2 shows that $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ for $p\leq d-1$ is not sufficient for infinitesimal Hilbertianity. This works towards answering Question 3 leaving only the interval of $p\in(d-1,d]$, where it is unknown whether $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ is sufficient to get infinitesimal Hilbertianity.